OpenStudy (anonymous):

I know this is probably easy, but the wording of it is messing with my brain.....

OpenStudy (anonymous):

After cutting a length of ribbon and giving 5/12 m of ribbon to Sam, Lily had 1/4 m of ribbon left. If she had 11/12 m of ribbon at the beginning, what was the length of ribbon Lily cut ? This is what I did... x - 5/12 = 1/4 x = 1/4 + 5/12 x = 3/12 + 5/12 x = 2/3 But I don't see what the 11/12 has to do with the problem ?

hartnn (hartnn):

maybe she had 1/4th of ribbon left means the initial 11/12 m plus the remaining left after giving

hartnn (hartnn):

**1/4m

OpenStudy (anonymous):

2/3 is the answer...right ?

hartnn (hartnn):

if remaining length 1/4 m equals total ribbon she has, then no...

OpenStudy (anonymous):

This is my sisters 6th grade homework. She asked me to help....the wording confused me. So how do I set this problem up ?

OpenStudy (anonymous):

11/12 - 5/12 - x = 1/4 ??

OpenStudy (anonymous):

11/12 - x - 5/12 = 1/4......= remaining ??

hartnn (hartnn):